OpenStudy (anonymous):

Please Someone Help ASAP. I WILL GIVE MEDALS 1. A spinner numbered 1 through 10 is spun 3 times. What is the probability of spinning a 5, then an even number, and then another even number? Write your answer as a decimal. Do not round. 2. A bag contains 3 red, 5 brown, and 4 yellow potatoes. What is the probability of randomly choosing 2 yellow potatoes without replacement? Write your answer as a decimal. Round to the nearest hundredth if needed.

OpenStudy (deadpoolroxz):

2:10

OpenStudy (deadpoolroxz):

for number 2

OpenStudy (anonymous):

1. P(5) = 1/10 P(even) =5/10 P(even2) = 5/10 total P = P(5) *P(even) * P(even2)